Can you step by step solve for x?
log (base 6) 1/216 = x

Answer by lwsshak3(11628)

Can you step by step solve for x?
log (base 6) 1/216 = x
..
log6(1/216)=x
convert to exponential form:
The base raised to the logarithm of the number=number. In this case the base=6,the logarithm of the number=x and the number=1/216
6^x=1/216=1/6^3=6^-3
exponents of the same base are equal to each other in an equation
x=-3